OpenStudy (smarty):

Find the determinant of the following matrix. [-4 2] [5 -2] A. 18 B.1 C.2 D. -2

OpenStudy (anonymous):

\[\det \left[\begin{matrix}-4 & 2 \\ 5 & -2\end{matrix}\right]=(-4)\times(-2)-2\times5=8-10=-2\] Thus, D. ;) Cheers

OpenStudy (unklerhaukus):

\[\left|\begin{array}{ccc} -4 & 2 \\ 5 & -2 \end{array}\right|\]\[=(-4\times-2)-(2\times5)\]\[=8-10\]\[=-2\]

OpenStudy (smarty):

Thanks to both!